Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 158     Entry time: 07 Sep 2020, 14:37
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 07.09.2020 
General Announcement:
- XO thanks instruments and support groups for using the E-logs while reporting activities. Please place names and groups in the entries to make a follow-up easier.

CTRL:
- SA3 vacuum loop is recovered.
- Today afternoon we want to move SA1 gui's to a more powerful server. If someone has an issue please contact CTRL.

DA:
- Not present.

EEE:
- Tests for MID to take place this afternoon.
- Cabling activities are ongoing.

VAC:
- Nothing to report.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- The video conference system needs to be tested in the control hutches. ITDM wants to set procedures with WOCO and also mounting it in experimental hutches. 
Instruments that are willing to help with the test can contact ITDM. The idea is to test in a Linux machine and test with remote users. SQS and HED are initially 
interested. 

XRO:
- Not present due to M3/SA2 re-alignment.

LAS:
- SA2 updates must be done by the end of today.
- SA1 and SA3 must be operational by the end of the day.

UNSYS:
- Requested ZZ to go to tunnel to fix communication issues with SA2/cell31. PRC arranged with RC to take place this morning.

SPB:
- Preparation for a couple of experiments next weeks.

FXE:
- Nothing to report.

MID:
- All fine at MID: checks with EEE to be done this afternoon and on Wednesday.

HED:
- Preparing M3 and other internal activities for user run next week.

SCS:
- Users will have beam time this week.
- SCS asks on the time window: Andreas G. says it is typically 400us but one can find in the delivery tables. 
If instruments request for more, please communicate to XO and OB will discuss it together with RCs.

SQS:
- Preparing next beam times.
- Shutter remote control: waiting for an update.
- Absorber in SA3 is currently in an unknown state. Nerea J. will check with Maurizio V.
ELOG V3.1.4-7c3fd00